A fraction has a reduced ratio of 2:5. When 5 is added to both the numerator and the denominator, the new reduced ratio is 1:2. Find the original fraction.
Answer by richwmiller(17219)

(2x+5)/(5x+5)=x/2x
x=5
check
10/25=2/5
ok
15/30=1/2
ok
original fraction is 10/25